How much water will overflow in a steel pipes hot water heating system containing 75 cubic foot of water? The Hot water heating system is filled with water at 40 °C and then heated to 90 °C. The coefficient of volume expansion for water is 0.00037 per C° or 0.0002 per F°.
1
Expert's answer
2020-11-25T07:12:15-0500

The volume increase "\Delta V" is given by the law of volume termal expansion:


"\Delta V = \alpha_VV\Delta T"

Here "\alpha_V = 0.00037\space \degree C^{-1}" is the coefficient of volume expansion for water, "V = 75 \space ft^3" is the initial volume, and "\Delta T = 90\degree C - 40\degree C = 50\degree C" is the increase in temperature. Thus, obtain:


"\Delta V = 0.00037\cdot 75\cdot 50 = 1.3875\space ft^3"

Answer. The overflow is "1.3875\space ft^3".